OpenStudy (anonymous):

A store is having a sale on 200 DVD movies. Some of the movies are being sold for $10 and some of them are being sold for $12. If the store made $2200 after selling all 200 DVDs, how many $12 DVDs were sold?

OpenStudy (anonymous):

$10x+$12y= $2200 x+y=200 multiple the bottom equation by -10 So you now $10x+$12y= $2200 -10x-10y=-2000 (the x's cancel out) nad you are left with 2y=200 y=100 so 100 $12 DVD's were sold

OpenStudy (anonymous):

thanks!